(b) Obtain the inverse Laplace transform for the following function: 5 +65 +9 (s- 2Xs+47/s-1)

differential equations
Answer #1

32+67+ 9 А B (3-2) (3+4) (1-1) s-2 1+4 82 +63 +9 $? (A +B+C) to (2A+ 38-31) -8A –48+20 A+B+C = 1 2A + 38-30 = 6 - PA - 4B +20